OpenStudy (anonymous):

3|-4x+75|+9=144 HELP

myininaya (myininaya):

3|-4x+75|=199-4 |-4x+75|=195/3 -4x+75=195/3 or -4x+75=-195/3

OpenStudy (anonymous):

Just solve for the absolute value part first, then break it into two different equations.

OpenStudy (anonymous):

Then solve each equation and the solution set is the union of the solutions for each equation.

OpenStudy (anonymous):

but the answer is suppose to be 30.. http://www.wolframalpha.com/input/?i=3 |-4x%2B75|%2B9%3D144 i don't know how they did it though

OpenStudy (anonymous):

LOL ^ LOVE YOU, but that is not the solution... it should be 15/2

myininaya (myininaya):

x=-5/2 and 70/2 is what I got let me check

myininaya (myininaya):

omg i cant read numbers i did it for 199 not 144 lol

OpenStudy (anonymous):

yeah 15/2 which is... 30

OpenStudy (anonymous):

loll

OpenStudy (anonymous):

how did you solve it though? can you take a picture of your work so i know how you do it?

OpenStudy (anonymous):

\[3|-4x+75| = 144-9\] \[|-4x + 75| = 135/3\] \(-4x + 75 = 135/3 \) or \(-4x + 75 = -135/3 \) \(-4x = 135/3 - 225/3\) or \(-4x = -135/3 -225/3\) \(-4x = -90/3\) or \(-4x = -360/3\) \(x = 30/4\) or \(x = 120/4\) \(x = 7/2\) or \(x = 30\)

OpenStudy (anonymous):

Hrm. I missed something on the left equation.

myininaya (myininaya):

x=15/2 or x=30

myininaya (myininaya):

you reduced wrong

OpenStudy (anonymous):

Yeah I see.

OpenStudy (anonymous):

\[x = 30/4 = 15/2 \ne 7/2\]

myininaya (myininaya):

at least when you read 4s you didn't write 9s and when you read 9s you didn't write 4s

OpenStudy (anonymous):

True ;p